# Visual Whole-Body Control for Legged Loco-Manipulation

### v2

@@ IV-B Real-World Experiments
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC as tested in the simulator directly into the real world. We In the real world, we test how many times the robot can pick up the assigned object in 5 trials (i.e., the robot is allowed to continuously retry four times when it fails to grasp without resetting). We consider it a success when the object is picked more than 0.1m higher than the plane where it is placed and a failure when the robot tries more than five times or the object falls from its surface. Each object is evaluated over three different heights, as stated before, and five resets are permitted for each height. In this part, we compare VBC with a similar high-level policy trained with stationary legs. To achieve the mobile manipulation task, we asked an experienced person to teleoperate with the robot using the default controller to get to an appropriate distance from the objects, and the trained policy would pick up the objects.
Figure 12: Visualization of robot behavior and corresponding commands during the test time.
Pickup performance. We collect all results and conclude the averaged performance over all objects on each height setting and show in Fig. 11. Obviously, VBC surpass the baseline method on all settings. To highlight, the baseline methods fail at 0.0m and 0.3m, as the default controller keeps a fixed robot height, similar to the floating base baseline. Even on the 0.5m setting, where the teleoperated baseline without whole-body behavior is equivalent to a static robotics arm, VBC is generally better. Besides, VBC shows great generalization ability on unseen shapes, thanks to the training strategy and observation design in the training pipeline.
@@
V Conclusions and Limitations
In this paper, we proposed a fully autonomous mobile manipulation system based on quadruped robots and a hierarchical training pipeline, Visual Whole-Body Control (VBC). VBC consists of a low-level goal-reaching policy and a high-level task-planning policy, which is trained through reinforcement learning (RL) and imitation learning (IL). Although we showed great success in picking various objects, we found several limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. Future works should consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.
@@ Acknowledgments
We would like to thank Unitree for their technical support. We thank Rizhao Qiu for providing the first version of the great TrackingSAM tool. We thank Chengzhe Jia for the great help on 3D modeling. We also thank Haichuan Che, Jun Wang, Shiqi Yang, and Jialong Li for their kind help in recording demos.
@@ II-A Low-Level Training
End-effector goal poses sampling. To train a robust low-level controller for goal-reaching, we should sample and track various goals (i.e., end-effector commands) as possible. To this end, we employ a height-invariant coordinate system originating at the robot’s base. This ensures the sampled trajectories are always on a sphere, guaranteeing the consistency of these samples; besides, it prevents the sampled goals from being affected by the quadruped robot’s height. In particular, we interpolate the end-effector goal command $\mathbf{p}^{\text{cmd}}$ between the current end-effector position $\mathbf{p}$ and a randomly selected end-effector position $\mathbf{p}^{\text{end}}$, at intervals of $T_{\text{traj}}$ seconds:
@@
$\mathbf{p}^{\text{end}}$ is resampled in a fixed timestep or if any $\mathbf{p}_{t}^{\text{cmd}}$ leads to self-collision or collision with the ground, similar to the linear and yaw velocity commands $v_{\text{lin}}^{\text{cmd}}$, $\omega_{yaw}^{\text{cmd}}$ that are uniformly sampled from a given range.
@@ II-B High-Level Training
Simulation We use the GPU-based Isaac Gym simulator [46]. As for the teacher, we train 10240 environments in parallel; regarding the student, we only utilize 240 parallel environments as there is a huge GPU memory cost for rendering visual observations. Experiments are trained on a single GPU (we use GTX4090 and GTX3090).

### v3

@@ IV-B Real-World Experiments
Evaluation principles and baselines. We deploy the trained visuomotor student policy of VBC as tested in the simulator directly into the real world. We In the real world, we test how many times the robot can pick up the assigned object in 5 trials (i.e., the robot is allowed to continuously retry four times when it fails to grasp without resetting), as we find that the robot emerges to have retrying behavior, as will be described below. We choose this as a measurement since we find that the emergent retrying behaviors improve the robustness, reliability, and success rate of finishing a job. We consider it a success when the object is picked more than 0.1m higher than the plane where it is placed and a failure when the robot tries more than five times or the object falls from its surface. Each object is evaluated over three different heights, as stated before, and five resets are permitted for each height. In this part, we compare VBC with a similar high-level policy trained with stationary legs. During the test of the real-world mobile manipulation task, we put the objects at an appropriate distance that the cameras mounted on the robot can observe, thus they can be manually annotated using our developed TrackingSAM tool. After the annotation, the robot is able to walk by and pick up the objects in a fully autonomous way.
Emergent retrying behaviors. It is worth noting that during simulation training, the robot emerges to learn a retrying behavior. In other words, when the robot fails to grasp an object, it will automatically retry to grasp without any human intervention. This shows an advantage and a clear difference compared to model-based methods.
Pickup performance. We collect all results and conclude the averaged performance over all objects on each height setting and show in Fig. 11. It is worth noting that our method allows for autonomous retrying behaviors when one pick-up trial fails. Obviously, VBC surpass the baseline method on all settings. To highlight, the baseline methods fail at 0.0m and 0.3m, as the default controller keeps a fixed robot height, similar to the floating base baseline. Even on the 0.5m setting, where the teleoperated baseline without whole-body behavior is equivalent to a static robotics arm, VBC is generally better. Besides, VBC shows great generalization ability on unseen shapes, thanks to the training strategy and observation design in the training pipeline.
@@
V Conclusions, Limitations and Future Works
In this paper, we proposed a fully autonomous mobile manipulation system based on quadruped robots and a hierarchical training pipeline, Visual Whole-Body Control (VBC). VBC consists of a low-level goal-reaching policy and a high-level task-planning policy, which is trained through reinforcement learning (RL) and imitation learning (IL). Although we showed great success in picking various objects, we found several limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. In our current pipeline, before the robot autonomously picks up objects, we must manually specify the target object in both cameras mounted on the robot. Thanks to the generalizability of trackingSAM, we find that as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to another view for automated picking, and free the robot from being specified in the between of the range of the two cameras, which improves the usability of our system. Since the requirements of annotation and object visibility are totally decoupled from our sim-to-real framework, in the future we can further improve and replace these modules with automated annotation algorithms and a search/exploration module to further improve the automation of the whole system. Future works should also consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.
@@ Acknowledgments
We would like to thank Unitree for their technical support. We thank Chengzhe Jia for the great help on 3D modeling. We also thank Haichuan Che, Jun Wang, Shiqi Yang, and Jialong Li for their kind help in recording demos.
@@ II-A Low-Level Training
Low-level commands sampling. To train a robust low-level controller for goal-reaching, we should sample and track various goals (i.e., base velocity commands and end-effector commands) as possible. As for the base velocity commands, we uniformly sample the linear velocity from a range of $[-0.6,0.6]$, and the angular velocity from a range of $[1.0,1.0]$. Regarding the end-effector, we employ a height-invariant coordinate system originating at the robot’s base. Such a design ensures the sampled trajectories are always on a sphere, guaranteeing the consistency of these samples; besides, it prevents the sampled goals from being affected by the quadruped robot’s height. In particular, we interpolate the end-effector goal command $\mathbf{p}^{\text{cmd}}$ between the current end-effector position $\mathbf{p}$ and a randomly selected end-effector position $\mathbf{p}^{\text{end}}$, at intervals of $T_{\text{traj}}$ seconds:
@@
$\mathbf{p}^{\text{end}}$ is resampled in a fixed timestep or if any $\mathbf{p}_{t}^{\text{cmd}}$ leads to self-collision or collision with the ground, similar to the linear and yaw velocity commands $v_{\text{lin}}^{\text{cmd}}$, $\omega_{yaw}^{\text{cmd}}$ that are uniformly sampled from a given range. Fig. 13 shows a screenshot of the low-level training, where the robot is required to track the given velocity commands and the end-effector pose.
Figure 13: A screenshot of low-level training, where the robot is required to track the given velocity commands (linear and yaw) and the end-effector pose.
@@ II-B High-Level Training
Simulation We use the GPU-based Isaac Gym simulator [47]. As for the teacher, we train 10240 environments in parallel; regarding the student, we only utilize 240 parallel environments as there is a huge GPU memory cost for rendering visual observations. Experiments are trained on a single GPU (we use GTX4090 and GTX3090). On a GTX4090, The teacher policy takes around 36 hours for training, and the student policy consumes around 48 hours.
Additional Training Techniques We design several training techniques with our prior knowledge to help the high-level policy to learn a well-performed, stable, not only in the simulation, but can also be transferred into the real world. We briefly introduce them in the following:
  1. 1.
    Action delay. We add a one-step action delay to cover the inference and execution in the real world.
  2. 2.
    Command clip curriculum. We gradually clip the linear velocity command during training, as we find that a large velocity helps accelerate the learning convergence but a slow velocity highly stabilizes the real world performance.
  3. 3.
    Reward curriculum. The command penalty reward $r_{\text{cmd}}$ is designed to encourage the robot to slow its velocity when it is close to the object to be picked.
  4. 4.
    Randomly changes the object position. During training, we take a small probability ($10\%$) to randomly change the object position and pose. This helps the robot to learn the retrying behaviors when it finds it fails during one trial.
  5. 5.
    Forcing stop when closing gripper. We forcibly set the velocity command to 0 when the gripper command is closing. This contributes a lot to a stop-then-pick behavior and thus improves the performance.
These training techniques are adopted for both teacher policy and student policy training. When training the student policy, we additionally randomize the camera latency as the policy will never obtain the immediate images at the time it infers, but some pictures that are a few milliseconds before, similar to the real world. We randomize the value by measuring the camera latency of the device we used in our real-world experiments.